What is the slope of the line parallel to the line y + 6 = 2x + 5?
Answer by mananth(16949)

y + 6 = 2x + 5
y=-6+5+2x
y=2x-1
compare with equation y=mx+b
m= slope = 2
the parallel line will have same slope